# RAGGED: Towards Informed Design of Retrieval Augmented Generation Systems

### v1

@@ Abstract
Retrieval-augmented generation (RAG) greatly benefits language models (LMs) by providing additional context for tasks such as document-based question answering (DBQA). Despite its potential, the power of RAG is highly dependent on its configuration, raising the question: What is the optimal RAG configuration? To answer this, we introduce the RAGGED framework to analyze and optimize RAG systems. On a set of representative DBQA tasks, we study two classic sparse and dense retrievers, and four top-performing LMs in encoder-decoder and decoder-only architectures. Through RAGGED, we uncover that different models suit substantially varied RAG setups. While encoder-decoder models monotonically improve with more documents, we find decoder-only models can only effectively use documents, despite often having a longer context window. RAGGED offers further insights into LMs’ context utilization habits, where we find that encoder-decoder models rely more on contexts and are thus more sensitive to retrieval quality, while decoder-only models tend to rely on knowledge memorized during training. 111Code/data for the RAGGED framework are available at https://github.com/neulab/ragged
@@
Jennifer Hsia††thanks: Equal contribution. Afreen Shaikh Zhiruo Wang Graham Neubig Carnegie Mellon University {jhsia2,afreens,zhiruow,gneubig}@cs.cmu.edu
@@ 1 Introduction
Retrieval-augmented generation (RAG) (Chen et al., 2017a; Lewis et al., 2020) is a technique widely applied to enhance the performance of top-performing LMs on knowledge-intensive generation tasks like document-based question answering (Karpukhin et al., 2020). Given a question, the technique includes using a retriever model to obtain multiple relevant passages (i.e. paragraphs) across potentially different documents, then inputting these passages to a reader model as additional contexts for generating an answer.
@@
To provide more concrete suggestions of the best practices under various cases, we introduce an analysis framework, RAGGED,222For “retrieval augmented generation generalized evaluation device”. to test RAG combinations on a suite of representative document-based question answering (DBQA) tasks, including open-domain datasets like Natural Questions (Kwiatkowski et al., 2019) and HotpotQA (Yang et al., 2018), which respectively focus on single-hop and multi-hop questions, as well as BioASQ, which targets the specialized, biomedical domain. To ensure a comprehensive evaluation, we incorporate both classic sparse and dense retrievers, BM25 (Robertson et al., 2009) and ColBERT (Khattab and Zaharia, 2020), and four top-performing reader models in encoder-decoder and decoder-only architectures, from Flan (Chung et al., 2022; Tay et al., 2023) and LLaMa (Touvron et al., 2023b) families, respectively.
We begin by exploring ”How many contexts can readers benefit from?” (§5). Our analysis identifies the optimal context quantity for individual LMs. We find that encoder-decoder models can effectively utilize up to 30 passages within their 2-token limit, whereas decoder-only models’ performance declines beyond 5 passages, despite having twice the size of the context limit (4).
Given this intriguing difference, we investigate the models’ context utilization behaviors (§6) and ask ”How reliant are models on provided contexts?”. We find that decoder-only models, which memorize more during training, exhibit comparatively less reliance on additional, test-time contexts. In contrast, encoder-decoder models, which memorize less during training, are more faithful to the provided contexts. This suggests that providing passages for context-reliant encoder-decoder models is beneficial, whereas it is less so for memory-reliant decoder-only models.
Given that some models are reliant on context, we also examine the importance of context quality by asking “How does the retriever quality affect readers’ contextualization behavior?” (§7) Our analysis considers two aspects: a retriever’s ability to identify high-quality passages and a reader’s response to varying passage quality. While dense, neural retrievers perform better on open-domain questions, sparse, lexical retrievers readily achieve comparable accuracy on special domains, with much less computation. Neural retrievers’ advantage readily benefits encoder-decoder models, especially for single-hop questions. However, the benefits are much less pronounced for decoder-only models and multi-hop questions.
@@
In summary, we demonstrate how RAGGED enables us to derive actionable insights about the conditions under which state-of-the-art RAG components combine to excel. We introduce a reusable framework that can easily be adapted to analyze new RAG components, such as retriever and reader models, as they evolve. We release our full dataset and code, aiming to provide the community with a deeper understanding of the nuanced interplay between context quantity, quality, and model architecture in RAG systems.
@@ 2.1 Framework Overview
RAG system components. For example, we vary the choice of the retriever (e.g., BM25, ColBERT), the reader family, (e.g., LLaMa2, FlanT5), and the max input token length.
The number of retrieved passages, denoted as . We vary from from 1 to 100, though find the most insightful variations in behavior occur before .
Data slices of data to examine. For example, we examine slices where the top- retrieved passages include gold passages and where they do not include gold passages.
@@
We vary the number of passages and the choice of reader to see how different model architectures and context limits affect the effective number of context passages a reader can process. Here, we evaluate all instances instead of specific subsets. While some models monotonically improve with more provided passages, others may have an early peak and have limited ability to utilize more passages.
@@
We focus on varying the choice of reader and the slice of instances to examine how different readers use context when the context quality varies. We examine the impact of context quality by analyzing slices of data where the top-k retrieved passages include gold passages and when they do not. As a result of this analysis, one can better understand how sensitive a reader is to positive context and how robust it is against incorrectly retrieved passages. We also investigate when models benefit from context and when models may be harmed by context.
@@
We focus on varying the retriever and data domain, and observing how well retrievers can perform based on the nature of the questions and how sensitive readers are to the quality of the retrieved passages. As reader models may have varied sensitivity to retrieval quality, one could select more appropriate models given the question characteristics and retrieval performance.
@@ 2.2 Implementation Details
Instruction: Give simple short one phrase answers for the questions based on the context Context: [passage, passage, , passage] Question: [the question of the current example] Answer: .
We truncate the Context to make sure the latter question is still included in the prompt. While performing generation for with top- passages with requires demanding computation, we sample to represent the general trend. More model implementation details can be found in §A.
@@ 3 Experimenting with RAG Systems
Below, we explain the representative retrievers and readers we analyze in our experiments.
@@ 3.1 Retriever
For retrieval, we employ two contrasting approaches: (1) sparse retrieval, based on lexical information, and (2) dense retrieval, based on neural models.
@@
BM25, a probabilistic retrieval model formulated by Robertson et al. (2009), leverages TF-IDF (Chen et al., 2017b) principles to estimate passage relevance via term weighting and passage length normalization. BM25 uses term-matching and hence experiences fewer out-of-domain failures when facing special-domain vocabulary, such as legal or medical texts.
@@
One of the best-performing neural-based retrievers is ColBERT (Khattab and Zaharia, 2020), i.e., contextualized late interaction over BERT. The transformer-based, contextualized embedding of ColBERT makes it more proficient than BM25 at identifying semantic similarities between queries and passages beyond lexical matching.
@@ 3.2 Reader
We compare four top-performing, open-source reader models of varied architectures: encoder-decoder models from the Flan family, and decoder-only models from the LLaMa2 family.
Flan Models
@@
LLaMa Models
The Llama2 models are decoder models. For LLaMa models (Touvron et al., 2023a), we adopt the latest LLaMa models (Touvron et al., 2023b) in 7B and 70B parameters, both having a context length of 4k tokens. The key feature is that LLaMa models are trained with reinforcement learning human feedback (RLHF).
@@ 4 Datasets and Evaluation Metrics
4.1 Dataset
We adopt three DBQA datasets from various domains (Wikipedia, biomedical) and of various types (single-hop, multi-hop, list, yes/no).
@@ 4.1 Dataset
We choose the Natural Questions (NQ) dataset (Kwiatkowski et al., 2019) to examine model performance on the most generic open-domain, single-hop questions. NQ questions are real user-search queries on Google. We adopt the KILT version (Petroni et al., 2021) of the dataset, which for each example, provides at least one gold relevant passage and one short answer.
@@
We choose HotpotQA (Yang et al., 2018) which provides challenging multi-hop questions. Each question requires reasoning over at least two passages to answer. While maintaining the same Wikipedia domain with the NQ dataset, HotpotQA enables comparisons of model reasoning skills over multiple evidence pieces.
@@
We choose BioASQ’s Task 11B (Krithara et al., 2023) with biomedical questions as a representative of special-domain questions. Our evaluation dataset is a compilation of the BioASQ Task 11B training and golden enriched set. BioASQ also presents challenging question types, such as lists and yes/no questions.
For NQ and HotpotQA datasets in the open domain, we retrieve passages from the Wikipedia corpus provided by the KILT benchmark (Petroni et al., 2021), with a total of 111M passages. For BioASQ, we use the PubMed Annual Baseline Repository for 2023 (of Medicine, 2023), with a total of 58M passages, which are either titles or passages of PubMed papers.
@@ 4.2 Metrics
Below we describe the metric definition for each instance. Overall performance on a dataset is computed by averaging over all instances.
4.2.1 Retrieval
Following Petroni et al. (2021), we evaluate retrieval performance using the recall@k metric.
Recall@k
For a given example with a query and ground-truth passage(s), recall@k measures the ratio of ground-truth passages among the top- retrieved passages. This measure can readily support retrieval for multi-hop questions (i.e., in HotpotQA) as well, where instead of retrieving any one of the gold passages, all passages along the reasoning path are necessary.
4.2.2 Reader Metrics
We evaluate the reader predictions using exact match and unigram F metrics.
Exact Match
Exact match (EM) measures if the model-generated output is the same as at least one of the ground-truth answers (Richardson et al., 2013). This metric requires models to be precise, yet may be overly strict when models produce verbose answers. Given the extractive nature of the NQ dataset (i.e., answers are spans in supporting passages), we use EM to evaluate NQ.
Unigram F
Unigram F is a well-adopted metric for QA, which quantifies the overlap of unigrams in the generated and reference texts, computed through the harmonic mean of precision and recall at the unigram level (Petroni et al., 2021). For each query, we compute the F score of the reader output against all possible gold answers and report the highest score. This is suitable if we do not need the generated answer to match the exact formatting and wording of the ground truth answer. For example, “five seasons” and “5 seasons” would fail EM but can get a partial credit of 0.5 by F. We evaluate HotpotQA and BioASQ with F metric.
@@ 5 Are More Contexts Always Better?
In this section, we study how models perform with various numbers of passages (§5.2), and if these results relate to context limits (§5.3).
5.1 RAGGED Analysis Guidelines
@@
Experiment with different numbers of context passages provided to the reader model. Start from a small number first since many reader models’ performance peaks before their context limits are reached.
@@
Expect a non-linear relationship between the number of contexts and model performance. Initially, reader performance may improve with more contexts due to the increased availability of “signal”, or supporting information. However, as the number of context passages increases, the amount of “noise”, or irrelevant, information also increases. As a result, reader performance may plateau or even degrade.
@@
Fitting as many contexts as the reader’s context limit can hold does not always guarantee optimal downstream performance. Using these experiments, one can better understand a reader’s ability to sift for “signal” among “noise”, and thereby find the optimal range of context passages for your model. Staying within this range maximizes performance without incurring unnecessary computational costs.
5.2 Number of Provided Contexts
We analyze the effect of the number of retrieved passages on reader model performance by providing the top- passages retrieved by ColBERT and evaluating the EM scores of answers generated by four reader models.
Results in Figure 3 reveal substantially different patterns between Flan and LLaMa models. On the NQ dataset, EM scores of both LLaMa models peak early at before continuing to decline steadily. In contrast, Flan models peak later around , and only level off instead of declining. When conditioning on a total of 30 passages, LLaMa 7B and LLaMa 70B underperform Flan models by about 17 to 33 point increase in exact match.
Similarly on multi-hop questions, Flan models can still benefit from more than 10 passages, whereas LLaMa can only benefit from 1 or 2 passages. In contrast, on the BioASQ dataset, both models have a small peak around 2 passages and roughly maintain the results afterwards.
These findings imply that encoder-decoder models can more effectively process tens of passages. In contrast, decoder-only models can use at most 2–3 passages in all cases. This observation underscores the necessity to carefully select the number of provided passages for different models to optimize their performance.
5.3 Context Limit vs. Early Truncation
Despite LLaMa models having twice the context length of Flan models (4 versus 2), they do not achieve superior performance when there is a large number of contexts. Nonetheless, this evidence alone can not allow us to claim that LLaMa models are worse than Flan models at processing long contexts. LLaMa’s longer context window could be a confounding factor that exposes the model to more (noisy) passages. To make a fair comparison, we truncate LLaMa model inputs by 2 tokens as we do to the Flan models, then evaluate LLaMa’s truncated-context performance.
As shown by the dashed lines in Figure 3, early truncation can only prevent LLaMa from further degrading after 15 passages on Wikipedia-domain questions but it does not improve results with smaller . The best-performing LLaMa (70B, truncated) still scores lower than the smallest Flan model, despite an almost 7 times difference in their number of parameters (70B vs. 11B).
This outcome challenges the assumption that models trained with extended context limits are inherently better at processing more contexts. Contrary to expectations, our results show that the short-window encoder-decoder models are more efficient in leveraging long contexts.
6 Context Utilization Habits
We study how readers utilize contexts, by examining how capable models are at answering questions without context, but instead simply from their pre-trained knowledge and generalization capabilities (§6.2). We then study how the readers incorporate test-time contexts with varied qualities (§6.3, §6.4).
6.1 RAGGED Analysis Guidelines

### v2

@@ Abstract
Retrieval-augmented generation (RAG) can significantly improve the performance of language models (LMs) by providing additional context for tasks such as document-based question answering (DBQA). However, the effectiveness of RAG is highly dependent on its configuration. To systematically find the optimal configuration, we introduce RAGGED, a framework for analyzing RAG configurations across various DBQA tasks. Using the framework, we discover distinct LM behaviors in response to varying context quantities, context qualities, and retrievers. For instance, while some models are robust to noisy contexts, monotonically performing better with more contexts, others are more noise-sensitive and can effectively use only a few contexts before declining in performance. This framework also provides a deeper analysis of these differences by evaluating the LMs’ sensitivity to signal and noise under specific context quality conditions. Using RAGGED, researchers and practitioners can derive actionable insights about how to optimally configure their RAG systems for their specific question-answering tasks. 111Code/data for the RAGGED framework are available at https://github.com/neulab/ragged
@@
Jennifer Hsia††thanks: Equal contribution. Afreen Shaikh∗ Zhiruo Wang Graham Neubig Carnegie Mellon University {jhsia2,afreens,zhiruow,gneubig}@cs.cmu.edu
@@ 1 Introduction
Retrieval-augmented generation (RAG) (Chen et al., 2017; Lewis et al., 2020) is a technique widely applied to enhance the performance of top-performing LMs on knowledge-intensive generation tasks like document-based question answering (Karpukhin et al., 2020). Given a question, the technique includes using a retriever model to obtain multiple relevant passages (i.e. paragraphs) across potentially different documents, then inputting these passages to a reader model as additional contexts for generating an answer.
@@
To provide more concrete suggestions of the best practices under various cases, we introduce an analysis framework, RAGGED,222For “retrieval augmented generation generalized evaluation device”. to test RAG combinations on a suite of representative document-based question answering (DBQA) tasks, including open-domain datasets like Natural Questions (Kwiatkowski et al., 2019) and HotpotQA (Yang et al., 2018), which respectively focus on single-hop and multi-hop questions, as well as BioASQ, which targets the specialized, biomedical domain. To ensure a comprehensive evaluation, we incorporate both classic sparse and dense retrievers, BM25 (Robertson et al., 2009) and ColBERT (Santhanam et al., 2021), and top-performing reader model families GPT (Brown et al., 2020), Claude (Enis and Hopkins, 2024), Flan (Chung et al., 2022; Tay et al., 2023), LLaMa (Touvron et al., 2023b) families, respectively.
We begin by exploring “How many contexts can readers benefit from?” (§5). Our analysis identifies the optimal context quantity for different readers. We find that while some readers’ performance improves monotonically as context quantity increases, other models’ performance peaks early and starts deteriorating, sometimes even below their no-context performance.
To understand these differences in more detail, we ask "How robust are reader models to irrelevant context when gold passages are present versus absent?" (§6). We analyze the readers under these data slices, and find that even with sufficient information, all readers still get distracted by the irrelevant context and degrade in performance as the number of contexts increases. The difference is that some models’ performance drops below their no-context baseline, while others are more robust and remain above the baseline.
Although it is useful to know model behavior under the presence and absence of signal, a practitioner may not have such detailed information at deployment time. Hence, we also examine reader sensitivity to context quality at the coarser level of retriever choice, and ask “How does the retriever choice impact reader performance?” (§7) For Wikipedia-based questions, we find that using a neural retriever offers a big advantage over using a lexical retriever, but this advantage only translates well for single-hop questions and not for multi-hop ones. For special-domain questions, we find that using a neural retriever only offers a minimal advantage over using a lexical retriever, but this advantage is amplified in reader performance.
@@
In summary, we demonstrate how RAGGED enables us to derive actionable insights about the conditions under which state-of-the-art RAG components combine to excel. We introduce a reusable framework that can easily be used to analyze new retriever and reader models, as they evolve. We release our full dataset and code, aiming to provide the community with a deeper understanding of the nuanced interplay between context quantity, quality, and model architecture in RAG systems.
@@ 2.1 Framework Overview
RAG system components. For example, we vary the choice of the retriever (e.g., BM25, ColBERT), the reader family, (e.g., GPT, Claude, Flan, LLaMa), and the max input token length.
The number of retrieved passages provided as context, denoted as . We vary from 1 to 50, though we find that most insightful variations in behavior occur before .
Slices of data to examine based on the quality of retrieved passages. By quality, we mean the presence of gold passages in the top- retrieved passages.
@@
We observe how different model architectures and context limits respond to an increasing , number of context passages. We observe two kinds of behaviors from different model families: some models’ performance monotonically improves as increases while others’ performance peaks early and deteriorates steadily.
@@
We analyze how effectively readers use context when the context quality varies. Specifically, by context quality, we mean whether there is sufficient information, or signal, in the context to answer the question. Whether a passage is considered a signal or noise is determined by whether it is marked as gold by human annotators. Using these slices, one can examine how sensitive a reader is to “signal” and how robust it is to the “noise”.
@@
We vary the retriever choice and analyze how much that affects reader performance on questions of different domains (Wikipedia v. PubMed) and complexity (single-hop v. multi-hop).
@@ 2.2 Implementation Details
Instruction: Give simple short one phrase answers for the questions based on the context Context: [passage1, passage2, , passagek] Question: [the question of the current example] Answer: .
We sample to represent the reader behavior. More details can be found in §A.
@@ 3 Experimenting with RAG Systems
We introduce the retrievers and readers.
@@ 3.1 Retriever
We employ two approaches: (1) a sparse retriever, based on lexical information, and (2) a dense retriever based on neural embeddings.
@@
BM25 is a probabilistic retrieval model (Robertson et al., 2009) that estimates passage relevance via term weighting and passage length normalization. BM25 relies on term-matching, and thus is supposed to be proficient at identifying lexical similarity especially in special domains.
@@
One of the best-performing neural-based retrievers is ColBERT (Santhanam et al., 2021), i.e., contextualized late interaction over BERT. ColBERT uses contextualized embeddings instead of term-matching as in BM25, thus is supposed to be better at identifying semantic similarities between queries and passages.
@@ 3.2 Reader
We analyze closed-source models from the GPT and Claude families, and open-source models from the Flan and LLaMa families.
Flan
@@
LLaMa
We use 7B and 70B LLaMa2 models (Touvron et al., 2023a, b) and the 8B and 70B LLaMa3 models. The LLaMa2 models have a context length of 4k tokens while LLaMa3 models have double the context length at 8k tokens. The LLaMa3 models have three major differences: 1) They use grouped query attention, which groups query heads to understand similar information better, 2) Their vocabulary is four times larger than that of LLaMa2, 3) They are trained on a seven times larger dataset than LLaMa2’s.
GPT
We use GPT-3.5-turbo model (Brown et al., 2020). This model has a context length of 16k tokens, and is a closed source model, so further details about model size are unknown.
Claude
We use Claude Haiku, which is Anthropic’s fastest and most compact model (Enis and Hopkins, 2024). The context window of 200k tokens is the largest of all the models we compare in this paper, but the model size is unknown since the model is closed-source.
@@ 4 Datasets and Evaluation Metrics
4.1 Datasets
We adopt three DBQA datasets from various domains (Wikipedia, biomedical) and of various complexity (single-hop, multi-hop). Details about the corpus of passages used for retrieval are in Table 2.
@@ 4.1 Datasets
We choose Natural Questions (NQ) (Kwiatkowski et al., 2019), a Wikipedia-based dataset, to examine how models perform on generic, open-domain, single-hop questions. NQ questions are real user-search queries on Google. We adopt the KILT version (Petroni et al., 2021) of the dataset, which provides one short phrase answer and at least one gold passage for each question.
@@
We choose HotpotQA (Yang et al., 2018), a multi-hop, Wikipedia-based dataset to examine how effectively models can identify multiple signal passages and reason over them together. Each question requires reasoning over at least two passages to answer.
@@
We choose BioASQ’s Task 11B (Krithara et al., 2023), a PubMed-based dataset, with biomedical questions to examine how models perform on special-domain questions. Our evaluation dataset is a compilation of the BioASQ Task 11B training and golden enriched set.
@@ 4.2 Metrics
Retriever Metric
We evaluate retrieval performance using the recall@k metric, following Petroni et al. (2021). For a given query, recall@k measures the fraction of ground-truth passages that are among the top- retrieved passages.
Reader Metric
We use unigram F1, which quantifies the overlap of unigrams in the reader output and gold answer(Petroni et al., 2021). For each query, we compute the F1 score of the reader output against the list of gold answers and report the highest score. We opt for this metric instead of exact match or substring match since many language models may respond correctly semantically but still fail these other metrics because of inexact wording or the tendency to answer in complete sentences instead of the short phrase format of the gold answers.
@@ 5 Are More Contexts Always Better?
In this section, we study how models perform with various amounts of context passages (§5.1), and if these results relate to context limits (§5.4). Our RAGGED analysis guidelines are:
@@
Vary the number of context passages provided to the reader model, starting from to 50, instead of directly providing as many passages as the model’s context limit can fit. This allows us to determine the point at which the model reaches peak performance, as many models do so before hitting their context limit.
@@
Initially, reader performance may improve with more contexts due to the increased availability of “signal”, or helpful information. However, as the number of context passages increases, the amount of “noise”, or irrelevant information, also increases. As a result, the reader performance may plateau or start degrading, sometimes even beyond the no-context performance.
@@
Fitting as many retrieved contexts as possible under the reader’s context limit does not always guarantee optimal downstream performance. Using these experiments, one can find the optimal range of in which a reader can most effectively sift for “signal” among “noise”. This optimal range is often not a large , thus identifying this can help practitioners and researchers save time and cost.
5.1 Reader Trends with Scaling Contexts
Results in Figure 3 reveal two typical trends in reader performance as we increase the number of ColBERT-retrieved context passages from to 20.
In one trend, the reader model’s performance steadily improves with increasing and then plateaus. This trend is observed in models such as the Flan models and the GPT-3.5 model, which all peak at without any noticeable decline afterward. For these models, providing a larger number of contexts is beneficial. However, note that for special domains, the exception is that Flan models peak earlier , but still follow the trend of leveling off as increases.
In the second trend, the reader model’s performance peaks early (at ) then steadily declines without leveling off as increases. This pattern is seen in models such as the LLaMa models and the Claude model. For these models, it is safer to provide a smaller number of contexts.
5.2 Comparing Optimal Performances
We compare reader performances at , their optimal number of context passages. Regardless of the domain or the number of passage hops required to answer the question, GPT-3.5 consistently ranks near the top.
For Wikipedia-domain questions, Flan models consistently rank high, with the larger LLaMa models performing either on par or close behind, followed by Claude Haiku, and then the smaller LLaMa models.
For special-domain questions, Claude Haiku rises to the second-best rank despite being one of the lower-ranked models for Wikipedia-domain questions. However, it’s important to note that Claude Haiku performs best when no context is provided, indicating that its high ranking is due to its extensive pre-trained knowledge rather than its ability to effectively use context. These observations underscore the necessity to carefully select to optimize reader performance.
One potential reason that some models’ performance deteriorates as increases is that some retrieved passages are irrelevant and unhelpful, so the inclusion of more passages may also result in the inclusion of more “noise”. In 6, we dig deeper into how readers perform under data slices of different informational qualities.
5.3 RAG vs. No-context
We evaluate reader performance without context () and compare it with the standard setting with top- passages from ColBERT (Figure 3). We find that GPT generally performs better with context than without, except on NQ where its pre-trained knowledge already allows it to achieve high performance without context.
Claude Haiku, on the other hand, quickly deteriorates and consistently performs worse with context than without. As a result, although Claude Haiku ranks high in no-context performance, it ranks lower when context is provided.
Flan models consistently perform better with context than without context. Their ability to use context effectively allows them to make up for their lack of pretrained knowledge, evidenced by how they rank high in peak top-k performance but low in no-context performance.
LLaMa models perform better with context than without context only for smaller values (). Although LLaMa3’s no-context performance surpasses LLaMa2’s, likely due to its larger training set and vocabulary, LLaMa3 struggles to use contexts as effectively and thus falls behind LLaMa2 as increases.
5.4 The Effect of Context Limit
Not all retrieved passages are helpful "signals" for answering the question; some may be "noise" that distracts the reader. Hence as more passages are included, more “noise” can be included too. The ability to sift for signal among noise is what determines the success of a reader, not merely the size of the reader’s context limit. For example, Claude Haiku has the largest context limit among the readers, yet it struggles to effectively utilize a larger number of contexts and peaks at a (Figure 3).
Given that a large context limit may include more noise, we examine whether truncating the reader input can help exclude noise and improve reader performance. We test this hypothesis by comparing LLaMa2’s performance under its original context length of 4k and a truncated length of . As shown by the dashed lines in Figure 3, early truncation has minimal effects on improving LLaMa2’s performance on Wikipedia-based questions, indicating that truncating the input does not provide a significant advantage. These observations suggest that context limit is not a key factor in configuring a RAG system.
6 Reader Robustness in the Presence and Absence of Gold Passages
We study how the readers respond to noise under contexts with gold context passages (§6.1) and without gold context passages (§6.2). Our RAGGED analysis guidelines are: